# Generated Contents Enrichment

### v1

@@ Abstract
In this paper, we investigate a novel artificial intelligence generation task, termed as generated contents enrichment (GCE). Different from conventional artificial intelligence contents generation task that enriches the given textual description implicitly with limited semantics for generating visually real content, our proposed GCE strives to perform content enrichment explicitly on both the visual and textual domain, from which the enriched contents are visually real, structurally reasonable, and semantically abundant. Towards to solve GCE, we propose a deep end-to-end method that explicitly explores the semantics and inter-semantic relationships during the enrichment. Specifically, we first model the input description as a semantic graph, wherein each node represents an object and each edge corresponds to the inter-object relationship. We then adopt Graph Convolutional Networks on top of the input scene description to predict the enriching objects and their relationships with the input objects. Finally, the enriched graph is fed into an image synthesis model to carry out the visual contents generation. Our experiments conducted on the Visual Genome dataset exhibit promising and visually plausible results.
@@ I Introduction
Although GCE is essential for filling the semantic richness gap between the short description and the expected image, it is challenging to learn to solve it. Firstly, the enriching content should be semantically relevant to the ones in the description and form appropriate inter-semantic relationships. Secondly, the enriched scene graph, composited by the description and its enriched semantics, should be structurally real, which follows the distribution of scene graphs from real-world images. Thirdly, since the existing text-to-image generation methods cannot ensure preserving all the textual information in their generated image, the synthesized image from GCE should reflect the input description during the enrichment process.
Towards solving the mentioned challenges and achieving GCE, we propose an intriguing approach that, to some extent, mimics the human reasoning procedure for hallucinating the enriching content. Particularly, our introduced method is designed with an end-to-end trainable architecture consisting of three principal stages. As an initial step, we represent the input description as a scene graph, wherein each node resembles a semantic, i.e., an object, and each edge corresponds to an inter-object relationship. In the first stage, we enrich the scene graph iteratively. In each iteration, we append a single object by modeling it as an unknown node and its related inter-object relationships as unknown edges, which are then predicted employing a Graph Convolutional Network (GCN) by aggregating the information from the known nodes and edges; thus, the enriching contents are semantically relevant with the ones in the description. Then, the enriched scene graph is fed into a pair of Scene Graph Discriminators to guarantee its structural reality. In the second stage, we synthesize the enriched image from the enriched scene graph by adopting an image generator [3]. The third stage simultaneously feeds the enriched image and the input description into the Visual Scene Characteriser and Image-Text Aligner, respectively, thus the generated image is visually real and preserves the information of the description.
Therefore, our contribution is a novel end-to-end trainable framework appointed to handle the proposed GCE task as the first attempt, to our best knowledge. This is executed by demonstrating the input scene descriptions as scene graphs and then enriching the semantics by employing GCNs. Afterward, the final enrichments are produced by feeding the enriched scene graphs into an image generator. In contrast to traditional text-to-image generation approaches, which solely focus on visual authenticity, we enhanced the generation by incorporating explicit visual and textual content enrichments that are semantically and structurally coherent with the provided input description. We assess our method on the real-world images and scene descriptions with rich semantics from the Visual Genome dataset [1], and achieve encouraging quantitative and qualitative outcomes.
@@ II-D Semantic Expansion
In prior literature, [93] addresses image outpainting by performing semantic extension applied to a scene graph constructed from a provided input image. Compared to the GCE task, they attempt to solve a separate problem, leading to different objectives and assessment criteria. In other words, we introduce and confront GCE, aiming to enhance the image synthesis quality. This is performed by enriching the same scene described by an input text description by incorporating supplementary related semantics and content.
@@ III Method
Figure -1: High-Level Overview of our end-to-end Generated Contents Enrichment framework. In Stage 1, the input scene graph (SG) is fed to the Scene Graph Enricher ${G_{sg}}$ to produce an enriched SG. Besides, a pair of local and global discriminators in the SG Critic $D_{sg}$ are jointly trained to differentiate between original and enriched SGs. These discriminators aid the enrichment process in constructing realistic, structurally coherent, and semantically meaningful scenes. In Stage 2, the Image Synthesizer $G_{im}$ leverages the resultant enriched SG to generate an image. In Stage 3, essential visual and textual scene characteristics are extracted in the Visual Scene Characterizer $S_{cf}$ and the Image-Text Aligner $M_{im\_sg}$. These two components ensure that the enriched image appropriately reflects the original description’s inherent characteristics.
Fig. -1 illustrates a high-level representation of our end-to-end model accepting a scene graph (SG) $g$ as input to simultaneously enrich the content in both textual and visual domains. Our proposed approach for GCE consists of three stages performed iteratively.
Initially, Stage 1: Generative Adversarial SG Enrichment accomplishes content enrichment at the object level leveraging Graph Convolutional Networks (GCNs). The SG Enricher, functioning as our generator, employs GCNs and Multi-Layer Perceptrons (MLPs) to identify the enriching objects and their inter/intra-relationships. The SG Critic, as a pair of local and global discriminators, ensures the enriched parts are realistic, structurally coherent, and semantically meaningful both on their own and within the context of the entire graph that represents the scene as a whole. Eventually, the enriched scene graph advances to Stage2: Image Generation, where it serves as the basis for synthesizing an enriched image containing richer relevant content. Lastly, this enriched image, combined with the input SG, progresses to Stage 3: Visual Evaluation & Alignment. Here, we aim to guarantee that the final generated output is visually plausible and faithfully reflects the fundamental scene characteristics outlined in the input prompt. This process involves the employment of a scene classifier and a pair of image/text encoders.
@@ III-B Graph Convolutional Network
Our Graph Convolutional Network (GCN) is inspired by SG2IM [3] and shares similarities with Convolutional Neural Networks (CNNs) in various aspects. Though GCNs are primarily designed to process graphs as inputs, they have the capacity to produce an output graph with a structure similar to the input. This is analogous to CNNs employing appropriate paddings. Each GCN is composed of multiple graph convolution (GConv) layers utilized to aggregate the data from nodes and edges, enabling the propagation of information throughout the graph. In this framework, the $l^{th}$ GConv layer, denoted as $GConv^{(l)}$ performs the message-passing as:
@@ III-C Stage 1: Generative Adversarial SG Enrichment
First, Enriching Object GCN $G_{sg,obj}$ takes an input SG $g$ to append an enriching object to the scene. Following the addition of objects in the previous step, Enriching Edge Detector $G_{sg,edg}$ takes the new graph and the hidden object vectors and appends edges to the scene. Finally, Enriching Predicate GCN $G_{sg,pred}$ employs the resulting enriched SG from the previous step to detect the predicates associated with the newly added enriching edges. Together, these three steps progressively enhance the input SG by introducing new objects, edges, and predicates, ultimately enriching the scene’s content iteratively.
@@
$G_{sg,edg}$ is designed with the utilization of two MLPs $\phi_{s}$ and $\phi_{o}$ sharing the same architecture but with different sets of weights. These MLPs receive several inputs: the initial input node feature vectors $V_{O}^{(0)}$, the enriching object $\hat{o}$, and hidden vectors $V_{O}^{(L)}$ extracted from the final layer of $G_{sg,obj}$. The MLPs $\phi_{s}$ and $\phi_{o}$ are specifically designed to capture information from the objects and subjects, respectively, within the context of directed inter-object relationships. This process is clearly illustrated in Eq. 8 and further elucidated in the supplementary material in Fig. 2.
@@
In the Enriching Edge Detector $G_{sg,edg}$, each node vector is transformed into another space to determine the location of the new edges within the enriched graph. These transformed object vectors are then compared using the dot product to detect the enriching edges. The similarity of node features in the new space indicates a higher probability of an edge existing between the two corresponding nodes, as reflected in the entries of $\hat{M}$, similar to an adjacency matrix. Consequently, nodes denoted as $o_{e}$ are selected that form edges $\hat{e}=(o_{e},.,\hat{o})$ or $\hat{e}=(\hat{o},.,o_{e})$, where the notation ”.” signifies a placeholder for the unknown predicate. This selection process for nodes forming enriching edges by considering the similarity of node features in the transformed space is summarized as:
$\hat{e}=G_{sg,edg}(V_{O}^{(0)},v_{\hat{o}},V_{O}^{(L)}).$ (9)
@@
Scene Graph Critic $D_{sg}$, as depicted in Fig. -1, indeed consists of a pair of discriminators trained jointly with the rest of the network. The high-level architecture of this critic draws inspiration from discriminators employed in GAN-based image inpainting [94]. Some additional details regarding the architecture of GCNs are provided in the supplementary materials.
@@ III-F Objective Function
$G_{sg,edg}$ produces a matrix $\hat{M}$ which resembles an adjacency matrix. However, each entry $\hat{m}_{i,j}$ within this matrix signifies the probability of an edge existing between $o_{i}$ as a subject and $o_{j}$ as an object. The Enriching Edge Detector contributes to the objective function with Edge Loss,
@@
During the inference phase, an $unknown\_obj$ node along with its corresponding $unknown\_pred$ edges are introduced into the input graph before being passed to $G_{sg}$. The enrichment process can be iteratively performed multiple times to progressively add additional content. At each iteration of this process, images may be generated from the resulting enriched intermediate scene graphs.
@@ IV-B ChatGPT Comparison Experiments
In case the input is represented as raw text rather than modeling as scene graphs, Large Language Models (LLMs) may be employed to expand the scene descriptions [97, 98, 99, 100, 101, 102, 103]. However, LLMs cannot efficiently and effectively perform the GCE task. First, simultaneous training of LLMs and image generators demands substantial training data and entails significant computational expenses. Secondly, despite the LLMs’ capacity to enrich the textual information, they lack the ability to comprehend the visual implications of the augmented content and measure its adequacy for generating a single real-world image. Thirdly, the direct utilization of LLMs might result in redundant content generation for image synthesis, particularly given that even the state-of-the-art generation method [2] struggles with handling exceedingly complex prompts.
To illustrate the superiority of our proposed framework over LLMs, we conducted a comparison utilizing ChatGPT [98] for the content enrichment task. This is presented qualitatively in Figure IV-B and quantitatively in Table II. The input descriptions are provided to ChatGPT, which generates enriched descriptions that are then used to synthesize images utilizing Stable Diffusion. We limited the word count for ChatGPT output as originally it generates very long texts unsuitable for inputs to SOTA image generators. These three different styles of prompts were used:
@@
Input Text
Enriched Text
@@
(Directly by Input Text)
@@
Figure 0: Qualitative Results. Samples of scene descriptions from the Visual Genome test split have been enriched, along with their synthesized images featuring richer content. The simple image is generated directly from the input text, and the enriched images are produced from their corresponding enriched text.
@@ IV-C Qualitative Results
Fig. IV-B showcases input descriptions and their enrichments alongside their corresponding synthesized images employing ChatGPT Direct and our proposed model. These input scene graphs are selected from the VG test split. In cases where the original scene graph contains a large number of objects, a subgraph with its associated edges is randomly selected and used as the initial input for our experiments. This subgraph strategy is primarily driven by the limitations of image generators when synthesizing images with lots of objects. Thus, if the initial input description already contains a substantial number of objects, the image synthesizer cannot accurately depict the enriched scene with even more objects. However, it is essential to note that there are no constraints regarding the number of semantics in the input description. Our proposed method carries out semantic enrichment iteratively, focusing on enriching one object at a time in each step, regardless of the number of existing objects in the scene.
@@ IV-D1 Inpainting Inspired Metrics
As so many terms are encompassed within our loss function, the contribution of each term plays an essential role in the overall enrichment task. For this reason, we conducted comprehensive hyperparameter tuning, and detailed results and discussions can be found in the supplementary materials. This includes the selected hyperparameters corresponding to the models in Table I, providing our quantitative evaluation methods for some of the chosen instances.

### v2

@@ Abstract
In this paper, we investigate a novel artificial intelligence generation task, termed as generated contents enrichment (GCE). Different from conventional artificial intelligence contents generation task that enriches the given textual description implicitly with limited semantics for generating visually real content, our proposed GCE strives to perform content enrichment explicitly on both the visual and textual domain, from which the enriched contents are visually real, structurally reasonable, and semantically abundant. Towards to solve GCE, we propose a deep end-to-end method that explicitly explores the semantics and inter-semantic relationships during the enrichment. Specifically, we first model the input description as a semantic graph, wherein each node represents an object and each edge corresponds to the inter-object relationship. We then adopt Graph Convolutional Networks on top of the input scene description to predict the enriching objects and their relationships with the input objects. Finally, the enriched description is fed into an image synthesis model to carry out the visual contents generation. Our experiments conducted on the Visual Genome dataset exhibit promising and visually plausible results.
@@ I Introduction
Although GCE is essential for filling the semantic richness gap between the short description and the expected image, it is challenging to learn to solve it. Firstly, the enriching content should be semantically relevant to the ones in the description and form appropriate inter-semantic relationships. Secondly, the enriched scene graph, composited by the description and its enriched semantics, should be structurally real, which follows the distribution of scene graphs from real-world images.
Towards solving the mentioned challenges and achieving GCE, we propose an intriguing approach that, to some extent, mimics the human reasoning procedure for hallucinating the enriching content. Particularly, our introduced method is designed with an end-to-end trainable architecture consisting of three principal stages. As an initial step, we represent the input description as a scene graph, wherein each node resembles a semantic, i.e., an object, and each edge corresponds to an inter-object relationship. In the first stage, we enrich the scene graph iteratively. In each iteration, we append a single object by modeling it as an unknown node and its related inter-object relationships as unknown edges, which are then predicted employing a Graph Convolutional Network (GCN) by aggregating the information from the known nodes and edges; thus, the enriching contents are semantically relevant with the ones in the description. Then, the enriched scene graph is fed into a pair of Scene Graph Discriminators to guarantee its structural reality. In the second stage, we synthesize the enriched image from the enriched scene description by adopting an image generator [3, 2]. The third stage simultaneously feeds the enriched image and the input description into the Visual Scene Characteriser and Image-Text Aligner, respectively, thus the generated image is visually real and preserves the information of the description.
Therefore, our contribution is a novel end-to-end trainable framework appointed to handle the proposed GCE task as the first attempt, to our best knowledge. This is executed by demonstrating the input scene descriptions as scene graphs and then enriching the semantics by employing GCNs. Afterward, the final enrichments are produced by feeding the enriched scene descriptions into an image generator. In contrast to traditional text-to-image generation approaches, which solely focus on visual authenticity, we enhanced the generation by incorporating explicit visual and textual content enrichments that are semantically and structurally coherent with the provided input description. We assess our method on the real-world images and scene descriptions with rich semantics from the Visual Genome dataset [1], and achieve encouraging quantitative and qualitative outcomes.
@@ II-D Semantic Expansion
The task of visual comprehension is addressed by generating an enriched and expressive scene graph as output from a given input image [93, 94]. This is performed by incorporating common-sense knowledge from the input images. Additionally, [95] focuses on the task of image outpainting, where an input image is expanded beyond its original boundaries to output an extrapolated version of the input. Their approach involves constructing and extending a scene graph based on the input image.
In contrast, our work introduces and tackles the task of GCE, aiming to enhance the image synthesis quality. While previous studies have engaged in semantic expansion, their objectives, input-output structures, and evaluation criteria differ significantly from ours. Specifically, the tasks addressed in the literature involve generating scene graphs from images, whereas our GCE approach involves transforming scene descriptions into enriched descriptions and generating their corresponding enriched images.
@@ III Method
Figure -1: High-Level Overview of our end-to-end Generated Contents Enrichment framework during the training phase. In Stage 1, the input scene graph (SG) is fed to the Scene Graph Enricher ${G_{sg}}$ to produce an enriched SG. Besides, a pair of local and global discriminators in the SG Critic $D_{sg}$ are jointly trained to differentiate between original and enriched SGs. These discriminators aid the enrichment process in constructing realistic, structurally coherent, and semantically meaningful scenes. In Stage 2, the Image Synthesizer $G_{im}$ leverages the resultant enriched SG to generate an image. In Stage 3, essential visual and textual scene characteristics are extracted in the Visual Scene Characterizer $S_{cf}$ and the Image-Text Aligner $M_{im\_sg}$. These two components ensure that the enriched image appropriately reflects the original description’s inherent characteristics.
Fig. -1 illustrates a high-level representation of our end-to-end model during the training phase. The model accepts a scene graph (SG) $g$ as input to simultaneously enrich the content in both textual and visual domains. Our proposed approach for GCE consists of three stages performed iteratively.
Initially, Stage 1: Generative Adversarial SG Enrichment accomplishes content enrichment at the object level leveraging GCNs. The SG Enricher, functioning as our generator, employs GCNs and Multi-Layer Perceptrons (MLPs) to identify the enriching objects and their inter/intra-relationships. The SG Critic, as a pair of local and global discriminators, ensures the enriched parts are realistic, structurally coherent, and semantically meaningful both on their own and within the context of the entire graph that represents the scene as a whole. Eventually, the enriched scene graph advances to Stage2: Image Generation, where it serves as the basis for synthesizing an enriched image containing richer relevant content. Lastly, this enriched image, combined with the input SG, progresses to Stage 3: Visual Evaluation & Alignment. Here, we aim to guarantee that the final generated output is visually plausible and faithfully reflects the fundamental scene characteristics outlined in the input prompt. This process involves the employment of a scene classifier and a pair of image/text encoders.
@@ III-B Graph Convolutional Network
Our Graph Convolutional Network (GCN) shares similarities with Convolutional Neural Networks (CNNs) in various aspects. Though GCNs are primarily designed to process graphs as inputs, they have the capacity to produce an output graph with a structure similar to the input. This is analogous to CNNs employing appropriate paddings. Each GCN is composed of multiple graph convolution (GConv) layers utilized to aggregate the data from nodes and edges, enabling the propagation of information throughout the graph. In this framework, the $l^{th}$ GConv layer, denoted as $GConv^{(l)}$ performs the message-passing as:
@@ III-C Stage 1: Generative Adversarial SG Enrichment
First, Enriching Object GCN $G_{sg,obj}$ takes an input SG $g$ to append an enriching object to the scene. Following the addition of objects in the previous step, Enriching Edge Detector $G_{sg,edge}$ takes the new graph and the hidden object vectors and appends edges to the scene. Finally, Enriching Predicate GCN $G_{sg,pred}$ employs the resulting enriched SG from the previous step to detect the predicates associated with the newly added enriching edges. Together, these three steps progressively enhance the input SG by introducing new objects, edges, and predicates, ultimately enriching the scene’s content iteratively.
@@
$G_{sg,edge}$ is designed with the utilization of two MLPs $\phi_{s}$ and $\phi_{o}$ sharing the same architecture but with different sets of weights. These MLPs receive several inputs: the initial input node feature vectors $V_{O}^{(0)}$, the enriching object $\hat{o}$, and hidden vectors $V_{O}^{(L)}$ extracted from the final layer of $G_{sg,obj}$. The MLPs $\phi_{s}$ and $\phi_{o}$ are specifically designed to capture information from the objects and subjects, respectively, within the context of directed inter-object relationships. This process is clearly illustrated in Eq. 8 and further elucidated in the supplementary material in Fig. 2.
@@
In the Enriching Edge Detector $G_{sg,edge}$, each node vector is transformed into another space to determine the location of the new edges within the enriched graph. These transformed object vectors are then compared using the dot product to detect the enriching edges. The similarity of node features in the new space indicates a higher probability of an edge existing between the two corresponding nodes, as reflected in the entries of $\hat{M}$, similar to an adjacency matrix. Consequently, nodes denoted as $o_{e}$ are selected that form edges $\hat{e}=(o_{e},.,\hat{o})$ or $\hat{e}=(\hat{o},.,o_{e})$, where the notation ”.” signifies a placeholder for the unknown predicate. This selection process for nodes forming enriching edges by considering the similarity of node features in the transformed space is summarized as:
$\hat{e}=G_{sg,edge}(V_{O}^{(0)},v_{\hat{o}},V_{O}^{(L)}).$ (9)
@@
Scene Graph Critic $D_{sg}$, as depicted in Fig. -1, indeed consists of a pair of discriminators trained jointly with the rest of the network. Some additional details regarding the architecture of GCNs are provided in the supplementary materials.
@@ III-F Objective Function
$G_{sg,edge}$ produces a matrix $\hat{M}$ which resembles an adjacency matrix. However, each entry $\hat{m}_{i,j}$ within this matrix signifies the probability of an edge existing between $o_{i}$ as a subject and $o_{j}$ as an object. The Enriching Edge Detector contributes to the objective function with Edge Loss,
@@
During the inference phase, an $unknown\_obj$ node along with its corresponding $unknown\_pred$ edges are introduced into the input graph before being passed to $G_{sg}$. Among the $unknown\_pred$ edges from the $unknown\_obj$ to all the other objects, we preserve only the edge with the highest probability of existence, as predicted by $G_{sg,edge}$. The enrichment process can be iteratively performed multiple times to progressively add additional content. At each iteration of this process, one object and one predicate are added to the description, enabling the generation of images from the resulting enriched intermediate scene graphs.
@@ IV Experiments
Given the absence of existing work on graphs that directly addresses the specific task we tackle, our primary focus is to demonstrate the promise of the proposed approach. We also include a comparison with ChatGPT as a Large Language Model (LLM). Our objective is to illustrate the feasibility of enriching generated content rather than surpassing state-of-the-art text/graph-to-image models, GNNs, or scene classifiers. Other end-to-end trainable modules with similar functionality can be integrated into our approach to achieve potentially enhanced performance.
During the inference phase, we utilize Stable Diffusion [2], a more advanced image synthesizer compared to SG2IM [3], which is employed during the training phase. This allows for enhanced visualization of details without the computational constraints imposed during training.
@@ IV-B ChatGPT Comparison Experiments
In case the input is represented as raw text rather than modeling as scene graphs, LLMs may be employed to expand the scene descriptions [98, 99, 100, 101, 102, 103, 104]. However, LLMs cannot efficiently and effectively perform the GCE task. First, simultaneous training of LLMs and image generators demands substantial training data and entails significant computational expenses. Secondly, despite the LLMs’ capacity to enrich the textual information, they lack the ability to comprehend the visual implications of the augmented content and measure its adequacy for generating a single real-world image. Thirdly, the direct utilization of LLMs might result in redundant content generation for image synthesis, particularly given that even the state-of-the-art generation method [2] struggles with handling exceedingly complex prompts.
To illustrate the superiority of our proposed framework over LLMs, we conducted a comparison utilizing ChatGPT [99] for the content enrichment task. This is presented qualitatively in Fig. IV-B and quantitatively in Table II. The input descriptions are provided to ChatGPT, which generates enriched descriptions that are then used to synthesize images.
We limited the word count for ChatGPT output as originally it generates very long texts unsuitable for inputs to SOTA image generators. These three different styles of prompts were used:
@@
Input Description
Enriched Description
@@
(Directly by Input Description)
@@
Figure 0: Qualitative Results. Samples of scene graphs from the Visual Genome test split as the input descriptions are enriched, along with their synthesized images featuring richer content. The simple image is generated directly from the input description, and the enriched images are produced from their corresponding enriched descriptions. The input scene graphs (a) and the enriched scene graphs generated by our model (c) are represented as textual descriptions.
@@ IV-C Qualitative Results
Fig. IV-B showcases input descriptions and their enrichments alongside their corresponding synthesized images employing ChatGPT Direct and our proposed model. These input scene graphs are selected from the VG test split. The input scene graphs, along with the enriched scene graphs generated by our model, are represented as textual descriptions in Fig. IV-B (a) and (c).
In cases where the original scene graph contains a large number of objects, a subgraph with its associated edges is randomly selected and used as the initial input for our experiments. This subgraph strategy is primarily driven by the limitations of image generators when synthesizing images with lots of objects. Thus, if the initial input description already contains a substantial number of objects, the image synthesizer cannot accurately depict the enriched scene with even more objects. However, it is essential to note that there are no constraints regarding the number of semantics in the input description. Our proposed method carries out semantic enrichment iteratively, focusing on enriching one object at a time in each step, regardless of the number of existing objects in the scene.
@@ IV-D1 Inpainting Inspired Metrics
As so many terms are encompassed within our loss function, the contribution of each term plays an essential role in the overall enrichment task. For this reason, we conducted comprehensive hyperparameter tuning. This includes the selected hyperparameters corresponding to the models in Table I, providing our quantitative evaluation methods for some of the chosen instances. SG Enrich denotes our proposed method with optimized hyperparameters. Deep G refers to a version utilizing a deeper generator architecture. Best NoD represents the optimal scenario without the introduced discriminator, while 128 BS signifies the use of a significantly larger mini-batch size. Detailed results and discussions can be found in the supplementary materials.